# Planning-Driven Programming: A Large Language Model Programming Workflow

### v1

@@ 6 Case Study
Figure 5 illustrates example message fragments from LPW in the 120th problem of HumanEval using the GPT-3.5 backbone. LPW successfully generates the correct program, while Baseline, SP, SD, and LDB all fail. This problem requires to return a sorted array with the maximum k numbers. However, in the problem description (block (a)), the unspecified order in the output array introduces ambiguity, confusing other methods. LPW struggles at the initial solution plan (block (c)), while the issue is clarified in the [Revised Plan], during plan verification (block (d)). The visible test (block (b)) delineates the reverse order in the return array after sorting in descending order. The initial code with print statements (block (f)) fails on the visible test since the array is not reversed. Subsequently, its execution trace is compared with the plan verification (block (e)) to identify this bug, as described in the [Error Analysis] in block (g). The refined code, which first sorts the array in descending order and then reverses the first $k$ elements into ascending order, successfully addresses this problem.
7 Related Work
Program Synthesis. Program synthesis remains an open challenge of generating a program within a target domain-specific language (DSL) from given specifications. One prevalent approach involves searching the large space of possible programs. For example, generalized planning whose solution is formalized as a planning program with pointers (Segovia-Aguas et al., 2024; Lei et al., 2023) has demonstrated promising results in synthesizing program solutions for abstract visual reasoning tasks (Lei et al., 2024) when the DSL is carefully designed. However, hand-crafted DSLs often suffer from limited generalization capacity, and the huge search space diminishes its effectiveness. Recently, large language models trained on vast corpora have excelled in natural language processing (NLP) tasks and have been extended to code generation e.g., GPT-series (Achiam et al., 2023; OpenAI, 2024), Llama-series (Meta, 2024; Roziere et al., 2023; Touvron et al., 2023), and Claude-series (Anthropic, 2024). LPW and SLPW leverage the strengths of LLMs in NLP tasks to generate intended solutions in natural language. These text-based solutions demonstrate high-quality logical reasoning steps and satisfactory accuracy, thereby effectively aiding subsequent code generation.
Prompting Techniques. To imitate the logical chain in human brain when tackling reasoning tasks, prompting methods direct LLMs to decompose problems into solvable sub-problems (Jiang et al., 2023; Zhou et al., 2023; Lightman et al., 2024; Dhuliawala et al., 2023) and progressively infer the correct answer with intermediate outputs, as exemplified by chain-of-thought prompting (Wei et al., 2022; Kojima et al., 2022). Inspired by these studies, LPW and SLPW decompose a text-to-code problem into several sub-problems described by the solution plan and follow the chain-of-thought prompting idea to verify the solution plan against visible tests with step-by-step analysis. The generated plan and its verification provide step-by-step natural language instructions for code generation, aiding LLMs in both the initial code development and subsequent refinements.
Code Refinement. Accurate program solutions often require iterative refinements due to model limitations (Zhong et al., 2024; Chen et al., 2023b; Shinn et al., 2023). Various interactive approaches have been proposed to optimize debugging performance in LLMs, such as human feedback Chen et al. (2024); Le et al. (2022); Wu et al. (2023), trained models (Huang et al., 2023b; Le et al., 2022; Yasunaga & Liang, 2021), LLM-generated explanations (Chen et al., 2023b; Madaan et al., 2023; Shinn et al., 2023; Tang et al., 2023), and execution results (Zhong et al., 2024; Holt et al., 2024; Tian & Chen, 2023). Current state-of-the-art LLM debuggers, such as Self-Debugging and LDB, repair various seed programs to create the program solution. However, they encounter difficulties when the initial code substantially deviates from the original intent. Besides, without safeguarding, the refined code frequently diverges from the problem specifications. In contrast, LPW and SLPW develop initial code that adheres to the validated intended solution through plan verification, minimizing deviations from the problem description. The plan verification further guides code refinement, ensuring alignment with the problem specifications.
@@ 8 Conclusion
We propose LPW, a large language model programming workflow, for text-to-code generation tasks, which enables LLMs to accurately draft an initial program and effectively correct bugs. LPW uses various advanced code generation techniques and efficiently incorporates them into a two-phase development model. We further present SLPW, a sampling variant of LPW, where multiple initial programs are generated and then competitively refined as necessary. We evaluate LPW and SLPW on well-established text-to-code generation benchmarks across various LLMs. LPW significantly improves code generation accuracy compared to other existing approaches. SLPW achieves new state-of-the-art Pass@1 accuracy, with 98.2% on HumanEval, 84.8% on MBPP, 64.0% on APPS, and 35.3% on CodeContests benchmarks using GPT-4o as the backbone. These results highlight the effectiveness of our workflow in generating high-quality code and underscore the benefits of incorporating sampling and debugging. In the future, additional visible tests automatically generated by LLMs (Chen et al., 2023a) can be explored to improve the performance of LPW and SLPW.
Appendix A Appendix
Input: A problem description $Q$, large language mold backbone $\mathcal{M}$, visible tests $T_{v}$, maximum iterations for solution generation $I_{s}$, $k$ plan samples, $q$ output solution plans and their verifications.
Output: A set of plans with their verifications $S_{output}=\{(\Pi_{1},\mathcal{A}(\Pi_{1},T_{v})),\ldots,(\Pi_{q},\mathcal{A}(\Pi_{q},T_{v}))\}$.
1 // plan generation
2 $plans\leftarrow\mathcal{M}(Q,k,t=0.4)$ ;
3  // generate $k$ initial $plans=\{\Pi_{1},\ldots\Pi_{k}\}$ with temperature $t=0.4$
4 $S_{output}\leftarrow\emptyset$;
5
[…]
7  // set current iteration $i_{s}$ to 0
8 $\mathrm{InitialUCB(len}(plans))$ ;
9  // initialize UCB algorithm with $\mathrm{len}(plans)$ arms
10 while  $i_{s}<I_{s}$ do
11       $\Pi\leftarrow\mathrm{SelectArm}(plans)$;
12        // select an arm $\Pi$ in $plans$ leveraging UCB algorithm
13       // plan verification generation
14       $\mathcal{A}(\Pi,T_{v})\leftarrow\mathcal{M}(Q,\Pi,T_{v})$;
15       $n,\Pi^{\prime}\leftarrow\mathcal{A}(\Pi,T_{v})$ ;
16        // number of visible tests $n$ where the plan verification derives an accurate final output, revised plan $\Pi^{\prime}$
17
18      if $n=\mathrm{len}(T_{v})$ then
19             // verification intermediate output check
20             $z\leftarrow\mathcal{M}(Q,\mathcal{A}(\Pi,T_{v}),\Pi)$ ;
21              // number of visible tests $z$ where the plan verification contains correct intermediate outputs
22             if $z=\mathrm{len}(T_{v})$ then
23                   $S_{output}\leftarrow S_{output}\cup(\Pi,\mathcal{A}(\Pi,T_{v}))$ ;
24                    // save $(\Pi$, $\mathcal{A}(\Pi,T_{v}))$ if no error in $\mathcal{A}$
25                   Delete($plans,\Pi$) ;
26                    // delete arm $\Pi$ from $plans$ in UCB
27
28                  if $\mathrm{len}(S_{output})=q$ then
29                         break ;
30                          // return first $q$ solution plans and their verifications
31
32            else
33                   UpdateConfidence($\Pi,z$) ;
34                    // update the confidence interval of $\Pi$ with reward $z$
35
36      else
37             $plans(\Pi)\leftarrow\Pi^{\prime}$ ;
38              // replace $\Pi$ in $plans$ with revised plan $\Pi^{\prime}$
39             UpdateConfidence($\Pi,n$) ;
40              // update the confidence interval of $\Pi$ with reward $n$
41
42      $i_{s}\leftarrow i_{s}+1$;
43
44if $\mathrm{len}(S_{output})=0$ then
[…]
46        // return empty set if neither a valid plan nor its verification is created
47
48return $S_{output}$;
49
Algorithm 1 SLPW: Solution Generation Phase
Input: A problem description $Q$, solution generation output $S_{output}$, large language mold backbone $\mathcal{M}$, visible tests $T_{v}$, maximum iterations for code implementation $I_{c}$.
Output: Final code $f$.
1 $programs\leftarrow\emptyset$;
2 $i_{c}\leftarrow 0$ ;
3  // set current iteration $i_{c}$ to 0
4 // initial code generation
5 for $\mathrm{plan}$ $\Pi$, $\mathrm{verification}$ $\mathcal{A}(\Pi,T_{v})$ $\mathrm{in}$ $S_{output}$ do
6      $f\leftarrow\mathcal{M}(Q,\Pi,\mathcal{A}(\Pi,T_{v}))$;
7        // generate an initial program $f$ for each $\Pi$ and $\mathcal{A}(\Pi,T_{v})$
8       $solved\leftarrow\mathrm{Exe}(f,T_{v})$ ;
9        // return a boolean result $solved$ after executing $f$ on $T_{v}$
10       if $solved$ then
11             return $f$ ;
12              // return $f$ when $f$ passes $T_{v}$
13
14      $programs\leftarrow programs\cup(f,\mathcal{A}(\Pi,T_{v}))$ ;
15        // record ($f$, $\mathcal{A}$) for further refinements
16
[…]
18  // initialize UCB algorithm with len($programs$) arms
19
20while  $i_{c}<I_{c}$ do
21       $f,\mathcal{A}(\Pi,T_{v})\leftarrow\mathrm{SelectArm}(programs)$;
22        // select an arm ($f$, $\mathcal{A}$) using UCB algorithm
23       $f_{p}\leftarrow\mathcal{M}(Q,f,\mathcal{A}(\Pi,T_{v}))$ ;
24        // add print statements in $f$ resulting in $f_{p}$
25       $n,\bar{t_{v}},\mathcal{T}(f,\bar{t_{v}}),\leftarrow\mathrm{Exe}(f_{p},T_{v})$ ;
26        // the number of solved visible tests $n$, execution trace $\mathcal{T}$ on the first failed visible test $\bar{t_{v}}$ after executing $f_{p}$ on $T_{v}$
27       if $n=\mathrm{len}(T_{v})$ then
28             return $f$ ;
29              // return $f$ when $f$ passes $T_{v}$
30
31      else
32             // error analysis
33             $\mathcal{O}(\mathcal{A},\mathcal{T})\leftarrow\mathcal{M}(Q,f,\mathcal{A}(\Pi,\bar{t_{v}}),\mathcal{T}(f,\bar{t_{v}}))$ ;
34              // error analysis $\mathcal{O}$ by comparing $\mathcal{A}$ with $\mathcal{T}$
35             // code explanation
36             $\mathcal{E}(f)\leftarrow\mathcal{M}(Q,f)$ ;
37              // generate explanation $\mathcal{E}$ for program $f$
38             // code refinement
39             $f^{\prime}\leftarrow\mathcal{M}(Q,f,\mathcal{E}(f),\mathcal{O}(\mathcal{A},\mathcal{T}))$ ;
40              // generate refined program $f^{\prime}$
41             $programs((f,\mathcal{A}(\Pi,T_{v})))\leftarrow f^{\prime}$ ;
42              // replace $f$ in $programs$ with refined program $f^{\prime}$
43             UpdateConfidence($(f,\mathcal{A}(\Pi,T_{v})),n$) ;
44              // update confidence of $(f,\mathcal{A}(\Pi,T_{v}))$ with reward $n$
45
46      $i_{c}=i_{c}+1$
47return $\mathrm{None}$ ;
48  // return none when no $f$ passes $T_{v}$ after reaching $I_{c}$
49
Algorithm 2 SLPW: Code Implementation Phase
A.1 SLPW Pseudo-Code
Algorithms 1 and 2 present the pseudo-code for the solution generation phase and code implementation phase in SLPW. In Algorithm 1, SLPW initially generates $k$ plan samples, $plans$, on Line 2 with a temperature of $t=0.4$. For all other LLM queries, $t$ is set to 0 by default to improve reproducibility. Lines 6-24 repeatedly select a plan $\Pi$ from $plans$ (Line 7), conduct verification (Line 9), and check the intermediate step outputs in the verification (Line 13) to generate $q$ solution plans along with their verifications as the output (Lines 14-18) when the plan verification and its intermediate outputs are successfully validated. In the solution generation phase, SLPW utilizes the UCB algorithm to select a plan $\Pi$ for further processing. SLPW treats each plan $\Pi$ as an arm and updates the confidence interval of $\Pi$ on Line 23 leveraging the number of visible tests $n$ where the plan verification derives an accurate final output, when verification fails on a visible test on Line 10, resulting in $n$ being smaller than the number of $T_{v}$ (Line 11). Additionally, SLPW updates the confidence interval of $\Pi$ on Line 20 using the number of visible tests $z$ where the plan verification contains accurate intermediate outputs, when erroneous intermediate values in the verification are detected on Line 13, leaning to $z$ being smaller than the number of $T_{v}$ (Line 14). The revised plan $\Pi^{\prime}$ is empty (Line 10) when the plan verification confirms consistency between the derived outputs and the ground-truth outputs across all visible tests. We note that, unlike the standard multi-armed bandit problem where the distribution for each arm remains stable, SLPW replaces each plan $\Pi$ with the revised plan $\Pi^{\prime}$ (Line 22) before updating confidence (Line 23) if the verification fails on a visible test and a revised plan $\Pi^{\prime}$ is generated (Line 10). We hypothesize that $\Pi^{\prime}$ remains closely related to $\Pi$, as typically only a few lines are changed. Therefore, the performance of $\Pi$ can offer valuable guidance for $\Pi^{\prime}$.
Algorithm 2 summarizes the code implementation phase. During this phase, SLPW takes the output of the solution generation phase as input to generate a set of initial programs, $programs$, on Lines 4-9. Lines 11-26 repeatedly select a program $f$ (Line 12), add print statements to $f$ resulting in $f_{p}$ (Line 13), execute it on visible tests $T_{v}$ (Line 14), and return $f$ for further assessment on the hidden tests when $f_{p}$ solves visible tests $T_{v}$ (Lines 15-16). Otherwise, SLPW generates the refined program $f^{\prime}$ (Line 23) with the error analysis $\mathcal{O}$ (Line 19) and the code explanation $\mathcal{E}$ (Line 21) when $f_{p}$ fails on $T_{v}$ and $\bar{t_{v}}$ is the first failed visible test (Line 14). The code implementation phase follows the same confidence interval update strategy as the solution generation phase due to the same hypothesize that the refine program $f^{\prime}$ is closely related to $f$. It uses the UCB algorithm to competitively refine each program $f$ while replacing $f$ with the refined $f^{\prime}$ (Line 24) and updating the confidence interval of $f$ with the number of solved visible tests $n$ (Line 25).
[…]
      12       89.0       89.0       89.6       89.6       89.0
      20       89.0       89.6       90.2       90.2       89.6
[…]
Table 5: Pass@1 accuracy of SLPW on HumanEval with GPT-3.5 varies by iterations (Iters) in both the solution generation phase and the code implementation phase, as well as by the number of output plans along with verifications, $q$, in the solution generation phase.
A.2 Parameter Study
SLPW involves four hyper-parameters: (1) the number of iterations in the solution generation phase, (2) the number of iterations in the code implementation phase, (3) the number of plan samples $k$, and (4) the number of output plans along with verifications $q$ for further code implementation. We use the same iterations for the solution generation and code implementation phases to simplify analysis. To identify the optimal parameters for SLPW to achieve the best performance, we vary $q$ from 1 to 5, set the number of iterations to 12 and 20, and configure $k=2\times q$ to ensure that sufficient plan samples generate enough verifications for further code implementation. The results in Table 5 reveal that larger iterations and $q$ values generally improve performance on the HumanEval benchmark. However, an excessive number, $q=5$, has a detrimental effect on performance compared to $q=3$ and $q=4$ with the same number of iterations. For the same $q$ settings, increasing the number of iterations tends to improve performance but consumes additional token resources. A larger $q$ value results in a greater number of initially developed programs in the code implementation phase, thereby raising the probability of passing visible tests. However, it also increases the risk of failing hidden tests due to less specific consideration on how to handle test cases during the initial program generation. Compared to $q=3$, setting $q=5$ results in a 2.4% improvement in instances where the generated program solves only the visible tests but fails the hidden tests, out of a total of 164 problems, with 12 iterations.
@@ A.2 Parameter Study
Table 6: Pass@1 accuracy (Acc) for the variations of LPW and SLPW with the GPT-3.5 backbone: -V denotes the exclusion of plan verification in both the solution generation phase and the code implementation phase; and -E stands for the exclusion of code explanation in the code implementation phase. Other metrics remain consistent with those in Table 4.
A.3 Additional Ablation Study
Table 6 shows the performance of the variants of LPW and SLPW on the HumanEval and MBPP benchmarks using GPT-3.5 as the LLM backbone. The suffix -V stands for the LPW and SLPW variants that exclude the plan verification in both the solution generation and code implementation phases. The suffix -E denotes removing the code explanation in LPW and SLPW when generating the refined program in the code implementation phase. Compared to LPW-S and SLPW-S in Table 4, which refine the incorrect seed programs, LPW-V and SLPW-V leverage the solution plan to generate the initial program without verifying the plan’s correctness, then follow the same debugging processes as LPW-S and SLPW-S.
The results of LPW-V and SLPW-V in Table 6 demonstrate the significance of plan verification in both initial code generation and further refinements. This aligns with our hypothesis that plan verification serves as the intended solution for visible tests, increasing the confidence of LLMs in initial code generation and subsequent refinements. Compared with LPW-S and SLPW-S, LPW-V and SLPW-V consider the solution plan when drafting initial programs. However, the effect of the unverified plan is limited, as only the performance of LPW-V and SLPW-V on MBPP is improved compared with the results of LPW-S and SLPW-S in Table 4. This is consistent with the results in Table 1, where Self-Planning (SP) shows minimal improvement compared to Baseline when both involve a one-time effort without debugging. The code explanation facilitates LLMs in aligning text-based error analysis with code implementation when locating and refining incorrect program lines. It shows a greater impact in LPW compared to SLPW, as evidenced by the results of LPW-E and SLPW-E in Table 6.
@@ A.3 Additional Ablation Study
Table 7: The percentage of different failure reasons for LPW and SLPW on the HumanEval benchmark with GPT-3.5 as the backbone. Missing Conditions and Logic Errors arise from the same issues in the plan and plan verification. Differ from Intended Solution indicates the plan and plan verification are manually classified as correct, while the generated code deviates, resulting in failure. No Code represents the absence of valid plan and plan verification in the solution generation phase, leading to failure after reaching the maximum number of iterations. Others denotes error program solutions caused by various reasons that differ from the previously listed categories.

### v2

@@ 6 Case Study
Figure 7 illustrates example message fragments from LPW in the 120th problem of HumanEval using the GPT-3.5 backbone. LPW successfully generates the correct program, while all other approaches fail. This problem requires returning a sorted array with the maximum k numbers. However, in the problem description (block (a)), the unspecified order in the output array introduces ambiguity, confusing other methods. LPW struggles at the initial solution plan (block (c)), while the issue is addressed in the [Revised Plan], during plan verification (block (d)). The visible test (block (b)) delineates the reverse order in the return array after sorting in descending order. The initial code with print statements (block (f)) fails on the visible test since the array is not reversed. Subsequently, its execution trace is compared with the plan verification (block (e)) to identify this bug, as described in the [Error Analysis] in block (g). The refined code, which first sorts the array in descending order and then reverses the first k elements into ascending order, successfully solves this problem.
@@ 7 Conclusion
We introduce LPW, a large language model programming workflow, designed for text-to-code generation tasks. LPW effectively integrates various advanced code generation techniques within a two-phase development model. A key innovation of LPW is the incorporation of plan verification, which enables LLMs to accurately draft an initial program and effectively correct bugs. We evaluate LPW on well-established text-to-code generation benchmarks across various LLMs. LPW significantly improves code generation accuracy compared to other existing approaches and achieves new state-of-the-art Pass@1 accuracy, with 98.2% on HumanEval, 84.8% on MBPP, 59.3% on LiveCode, 62.6% on APPS, and 34.7% on CodeContests benchmarks using GPT-4o as the backbone. In the future, additional visible tests automatically generated by LLMs (Chen et al., 2023a) can be explored to improve the performance of LPW.
8 Limitations
Similar to other code generation approaches, LPW is constrained by the imperfect reasoning capacity of LLMs. Strengthening the reasoning capacity of LLMs remains an ongoing challenge. While the plan and plan verification have proven valuable across different benchmarks, they require substantial token consumption. In the future, reducing this consumption remains a critical area for improvement. Besides, although the LLM-generated plan and plan verification demonstrate promising accuracy on current tasks, the accuracy of the generated code still lags behind that of the plan and plan verification (Appendix E). Incorporating alternative solution representations (Zelikman et al., 2023) alongside natural language representations may assist LLMs in translating text-based solutions into program solutions more accurately.
Appendix
Appendix A Related Work
A.1 Program Synthesis
Program synthesis remains an open challenge of generating a program within a target domain-specific language (DSL) from given specifications. One prevalent approach involves searching the large space of possible programs. For example, generalized planning whose solution is formalized as a planning program with pointers (Segovia-Aguas et al., 2024; Lei et al., 2023) has demonstrated promising results in synthesizing program solutions for abstract visual reasoning tasks (Lei et al., 2024) when the DSL is carefully designed. However, hand-crafted DSLs often suffer from limited generalization capacity, and the huge search space diminishes its effectiveness. Recently, large language models trained on vast corpora have excelled in natural language processing (NLP) tasks and have been extended to code generation e.g., GPT-series (Achiam et al., 2023; OpenAI, 2024), Llama-series (Meta, 2024; Roziere et al., 2023; Touvron et al., 2023), and Claude-series (Anthropic, 2024). LPW leverages the strengths of LLMs in NLP tasks to generate intended solutions in natural language. These text-based solutions demonstrate high-quality logical reasoning steps and satisfactory accuracy, thereby effectively aiding subsequent code generation.
A.2 Prompting Techniques
To imitate the logical chain in human brain when tackling reasoning tasks, prompting methods direct LLMs to decompose problems into solvable sub-problems (Jiang et al., 2023; Zhou et al., 2023; Lightman et al., 2024; Dhuliawala et al., 2023) and progressively infer the correct answer with intermediate outputs, as exemplified by chain-of-thought prompting (Wei et al., 2022; Kojima et al., 2022). Inspired by these studies, LPW decomposes a text-to-code problem into several sub-problems described by the solution plan and follows the chain-of-thought prompting idea to verify the solution plan against visible tests with step-by-step analysis. The generated plan and its verification provide step-by-step natural language instructions for code generation, supporting LLMs in both the initial code development and subsequent refinements.
Figure 8: Pass@1 accuracy as a function of token consumption for LDB, Repeated Sampling, and LPW on the LiveCode benchmark with GPT-4o.
A.3 Code Refinement
Accurate program solutions often require iterative refinements due to model limitations (Zhong et al., 2024; Chen et al., 2023b; Shinn et al., 2023). Various interactive approaches have been proposed to optimize debugging performance in LLMs, such as human feedback (Chen et al., 2024; Le et al., 2022; Wu et al., 2023), trained models (Huang et al., 2023b; Le et al., 2022; Yasunaga and Liang, 2021), LLM-generated explanations (Chen et al., 2023b; Madaan et al., 2023; Shinn et al., 2023; Tang et al., 2023), execution results (Zhong et al., 2024; Holt et al., 2024; Tian and Chen, 2023), and multi-agent collaborations (Islam et al., 2024; Huang et al., 2023a; Qian et al., 2024; Dong et al., 2023b). Current state-of-the-art LLM debuggers, such as Self-Debugging and LDB, repair various seed programs to create program solutions. However, they encounter difficulties when the initial code substantially deviates from the original intent. Besides, without safeguarding, the refined code frequently diverges from the problem specifications. In contrast, LPW develops initial code that adheres to the validated intended solution through plan verification, minimizing deviations from the problem description. The plan verification further guides the code refinement, ensuring alignment with the problem specifications.
Figure 9: Pass@1 accuracy of LPW on the HumanEval benchmark using GPT-3.5 with 20 iterations in both the solution generation and code implementation phases.
Appendix B Parameter Study
We conduct an experiment involving 20 iterations for both the solution generation and code implementation phases in LPW. Figure 9 shows the variation in Pass@1 accuracy on the HumanEval benchmark using GPT-3.5. The increased number of iterations in the solution generation phase results in higher initial program accuracy, 81.1%, compared to 79.9% with 12 iterations before refinements. Subsequently, the accuracy steadily increases, reaching the highest value of 89.6% after 18 debugging turns, compared to a maximum of 89.0% with 12 iterations. Overall, a larger number of iterations improves performance in both initial code generation and subsequent refinements. However, the significant token consumption presents challenges for practical applications.
Appendix C Additional Ablation Study
@@
Table 5: Pass@1 accuracy (Acc) for the variant of LPW with the GPT-3.5 backbone. The suffix -E stands for the exclusion of code explanation in the code implementation phase. Other metrics remain consistent with those in Table 4.
@@
Table 6: The percentage of different failure reasons for LPW on the HumanEval benchmark with GPT-3.5 as the backbone. Missing Conditions arises from the same issues in the plan and plan verification. Differ from Intended Solution indicates the plan and plan verification are manually classified as correct, while the generated code deviates, resulting in failure. No Code represents the absence of valid plan and plan verification in the solution generation phase, leading to failure after reaching the maximum number of iterations. Others denotes error program solutions caused by various reasons that differ from the previously listed categories.
Table 5 shows the performance of the variant of LPW on the HumanEval and MBPP benchmarks using GPT-3.5 as the LLM backbone. The suffix -E denotes removing the code explanation when generating the refined program in the code implementation phase. The code explanation facilitates LLMs in aligning text-based error analysis with code implementation when locating and refining incorrect program lines. LPW-E demonstrates a decrease in Pass@1 accuracy on both the HumanEval and MBPP benchmarks.
Figure 10: Pass@1 accuracy of Baseline, Self-Planning (SP), MapCoder, Self-Debugging (+Expl) (SD), LDB, and LPW across different difficulty levels, Easy, Medium, and Hard on the HumanEval benchmark when leveraging GPT-3.5 as the LLM backbone.
Appendix D Analysis of Unsolved Problems for LPW using GPT-3.5
D.1 Performance Across Different Difficulty Levels
Figure 10 compares the Pass@1 accuracy of competing methods across different difficulty levels, Easy, Medium, and Hard, on the HumanEval benchmark using GPT-3.5. We apply the method described in Zhong et al. (2024) to generate the difficulty annotations in Figure 10 by querying GPT-3.5 with problem descriptions and canonical solutions. LPW displays convincing performance, exceeding 85% accuracy across all difficulty levels. For the Hard level, LPW achieves 85.7% accuracy, while competing approaches show a notable degradation, falling below 70%.